OpenStudy (anonymous):

1. find the perimeter of the rectangle. a) 11 in b) 22 in c) 28 in d) 56 in

pooja195 (pooja195):

Where is the rectangle?

OpenStudy (solomonzelman):

Perimeter of any shape is a sum of its sides.

pooja195 (pooja195):

See those lines? 

pooja195 (pooja195):

This means the the two sides are congruent

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Blue + blue = ? 7+7=?

OpenStudy (anonymous):

14

pooja195 (pooja195):

ok good red + red 4+4=?

OpenStudy (anonymous):

8

pooja195 (pooja195):

14+8=?

OpenStudy (anonymous):

22

OpenStudy (anonymous):

P=2(l+w) P=2(7+4) = 22 P=22

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

Good so our final answer is 22in.

OpenStudy (anonymous):

yes

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)